OpenStudy (anonymous):

how would i solve ln(√x+7) = 4

OpenStudy (anonymous):

\[\ln (\sqrt{x}+7)=4\]

OpenStudy (a_clan):

\[e ^{4}=\sqrt{x+7}\] \[e ^{8}=x+7\] \[x=e ^{8}-7\]

OpenStudy (anonymous):

thanks! how would i find the inverse of \[f \left( x \right)=\sqrt[3]{x+3}\]